# TRADES: Generating Realistic Market Simulations with Diffusion Models

### v2

@@ Abstract.
Financial markets are complex systems characterized by high statistical noise, nonlinearity, volatility, and constant evolution. Thus, modeling them is extremely hard. Here, we address the task of generating realistic and responsive Limit Order Book (LOB) market simulations, which are fundamental for calibrating and testing trading strategies, performing market impact experiments, and generating synthetic market data. Previous works lack realism, usefulness, and responsiveness of the generated simulations. To bridge this gap, we propose a novel TRAnsformer-based Denoising Diffusion Probabilistic Engine for LOB Simulations (TRADES). TRADES generates realistic order flows as time series conditioned on the state of the market, leveraging a transformer-based architecture that captures the temporal and spatial characteristics of high-frequency market data. There is a notable absence of quantitative metrics for evaluating generative market simulation models in the literature. To tackle this problem, we adapt the predictive score, a metric measured as an MAE, by training a stock price predictive model on synthetic data and testing it on real data. We compare TRADES with previous works on two stocks, reporting a ×$\times 3.27$ and ×$\times 3.48$ improvement over SoTA according to the predictive score, demonstrating that we generate useful synthetic market data for financial downstream tasks. Furthermore, we assess TRADES’s market simulation realism and responsiveness, showing that it effectively learns the conditional data distribution and successfully reacts to an experimental agent, giving sprout to possible calibrations and evaluations of trading strategies and market impact experiments. We developed DeepMarket, the first open-source Python framework for market simulation with deep learning. In our repository, we include a synthetic LOB dataset composed of TRADES’s generated simulations. We release the code at https://github.com/LeonardoBerti00/DeepMarket.
Diffusion Model, Market Simulation, Limit Order Book
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; August 03–07, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Neural networks
1. Introduction
@@
A realistic and responsive market simulation [^1] To abbreviate, throughout this paper, by “market simulation”, we refer to a limit order book market simulation for a single stock. has always been a dream in the finance world (Raberto and Cincotti, 2005; Levy et al., 2000; Mizuta, 2016; Maeda et al., 2020; Jacobs et al., 2004; Raberto et al., 2001). Recent years have witnessed a surge in interest towards deep learning-based market simulations (Li et al., 2020; Nagy et al., 2023; Cont et al., 2023; Coletta et al., 2021). An ideal market simulation should fulfill four key objectives: (1) enable the calibration and evaluation of algorithmic trading strategies including reinforcement learning models (Zhang et al., 2020); (2) facilitate counterfactual experiments to analyze (2.1) the impact of orders (Webster, 2023), and (2.2) the consequences of changing financial regulations and trading rules, such as price variation limits; (3) analyze market statistical behavior and stylized facts in a controlled environment; (4) generate useful granular market data to facilitate research on finance and foster collaboration.
@@
For these objectives, two key elements are paramount: the realism and the responsiveness of the simulation. Realism (Naeem et al., 2020) refers to the similarity between the generated probability distribution and the actual data distribution. Responsiveness captures how the simulated market reacts to an experimental agent’s actions. Furthermore, the generated data usefulness is crucial for achieving the last objective (4). Usefulness refers to the degree to which the generated market data can contribute to other related financial tasks, such as predicting the trends of stock prices (Prata et al., 2023; Sirignano, 2019; Tsantekidis et al., 2017b).
@@
Backtesting (Harvey and Liu, 2015) and Interactive Agent-Based Simulations (IABS) (Byrd et al., 2020) are two of the most used traditional market simulation methods. Backtesting assesses the effectiveness of trading strategies on historical market data. It is inherently non-responsive since there is no way to measure the market impact of the considered trading strategies, making the analysis partial. Minimizing market impact has been the focus of extensive research efforts over many years (Almgren, 2003), resulting in the development of sophisticated algorithms designed to mitigate the price effects of executing large orders, through timing, order, size, and venue selection. IABS, on the other hand, enables the creation of heterogeneous pools of simplified traders with different strategies, aiming to approximate the diversity of the real market. However, obtaining realistic multi-agent simulations is challenging, as individual-level historical data of market agents is private, which impedes the calibration of the trader agents, resulting in an oversimplification of real market dynamics.
@@
Recent advances in generative models, particularly Wasserstein GANs (Coletta et al., 2023, 2022, 2021; Li et al., 2020), have shown promise in generating plausible market orders for simulations. However, GANs are susceptible to mode collapse (Thanh-Tung and Tran, 2020) and training instability (Chu et al., 2020), leading to a lack of realism and usefulness in the generated data. These limitations hinder their ability to satisfy the market simulation objectives and their real-world applicability.
@@
To address these shortcomings, we present our novel TRAnsformer-based Denoising Diffusion Probabilistic Engine for LOB market Simulations (TRADES). TRADES generates realistic high-frequency market data, [^2] They refer both to orders and LOB snapshots (see details in Sec. 2.2 and Sec. 2). which are time series, conditioned on past observations. We demonstrate that TRADES surpasses state-of-the-art (SoTA) methods in generating realistic and responsive market simulations. Importantly, due to its ability to handle multivariate time series generation, TRADES is adaptable to other domains requiring conditioned sequence data generation. Furthermore, TRADES readily adapts to an experimental agent introduced into the simulation, facilitating counterfactual market impact experiments. In summary, our contributions are:
  1. (1)
@@
    Realistic and responsive market simulation method: We develop a Denoising Diffusion Probabilistic Engine for LOB Simulations (TRADES), exploiting a transformer-based neural network architecture. We remark that TRADES is easily adaptable to any multivariate time-series generation domain.
  2. (2)
@@
    Plug-and-play framework: We release DeepMarket, the first open-source Python framework for market simulation with deep learning. We also publish TRADES’s implementation and checkpoints to promote reproducibility and facilitate comparisons and further research.
  3. (3)
@@
  4. (4)
    New “world” agent for market simulations: We extend ABIDES (Byrd et al., 2020), an agent-based simulation environment, in our framework by introducing a new world agent class accompanied by a simulation configuration, which, given in input a trained generative model, creates limit order book market simulations. Our experimental framework does not limit the simulation to a single-world agent but enables the introduction of other trading agents, which interact among themselves and with the world agent. This defines a hybrid approach between deep learning-based and agent-based simulations.
  5. (5)
    First quantitative metric for market simulations: The literature shows a notable absence of quantitative metrics to evaluate the generated market simulations. Typically, the evaluation relies on plots and figures. We posit that a robust and quantitative assessment is essential for the comparative analysis of various methodologies. To this end, we adapt the predictive score introduced in (Yoon et al., 2019) to objectively and quantitatively evaluate the usefulness of the generated market data.
  6. (6)
    Extensive experiments assessing usefulness, realism, and responsiveness: We perform a suite of experiments to demonstrate that TRADES-generated market simulations abide by these three principles. We show how TRADES outperforms SoTA methods (Byrd et al., 2020; Coletta et al., 2022, 2021; Vyetrenko et al., 2020) according to the adopted predictive score and illustrate how TRADES follows the established stylized facts in financial markets (Vyetrenko et al., 2020).
2. Background
Here, we provide background information on multivariate time series generation and limit order book markets. Furthermore, since TRADES is an extension of the Denoising Diffusion Probabilistic Model (DDPM), we summarize it in the Appendix A.
2.1. Multivariate time series generation
@@
(1) $q(\mathbf{x})=\prod_{n=1}^{N}q(\mathbf{x}_{n}\>|\>\mathbf{x}_{1},...,\,\mathbf{x}_{n-1})$
@@
2.2. Limit Order Book
@@
In a Limit Order Book (LOB) market, traders can submit orders to buy or sell a certain quantity of an asset at a particular price. There are three main types of orders in a limit order market. (1) A market order is filled instantly at the best available price. (2) A limit order allows traders to decide the maximum (in the case of a buy) or the minimum (in the case of a sell) price at which they want to complete the transaction. A quantity is always associated with the price for both types of orders. (3) A cancel order [^5] Sometimes, events referring to these orders are defined as deletion. removes an active limit order. The Limit Order Book (LOB) is a data structure that stores and matches the active limit orders according to a set of rules. The LOB is accessible to all the market agents and is updated with each event, such as order insertion, modification, cancellation, and execution. The most used mechanism for matching orders is the Continuous Double Auction (CDA) (Bouchaud et al., 2018). In a CDA, orders are executed whenever a price overlaps between the best bid (the highest price a buyer is willing to pay) and the best ask (the lowest price a seller is willing to accept). This mechanism allows traders to trade continuously and competitively. The evolution over time of a LOB represents a multivariate temporal problem. We can classify the research on LOB data into four main types of studies, namely empirical studies analyzing the LOB dynamics (Cont, 2001; Bouchaud et al., 2002), price and volatility forecasting (Zhang et al., 2019; Sirignano, 2019), modeling the LOB dynamics (Cont, 2011; Gould et al., 2013) and LOB market simulation (Byrd et al., 2020; Coletta et al., 2021; Li et al., 2020).
3. Related Works
Diffusion models for time series generation. Diffusion models have been successfully applied to generate images (Austin et al., 2021; Karras et al., 2022), video (Mei and Patel, 2023), and text (Zhu et al., 2023). Recently, they have been exploited also for time series forecasting (Rasul et al., 2021; Li et al., 2022), imputation (Biloš et al., 2023; Tashiro et al., 2021), and generation (Lim et al., 2023). To the best of our knowledge, only Lim et al. (Lim et al., 2023) tackle time series generation using diffusion models. They present TSGM, which relies on an RNN-based encoder-decoder architecture with a conditional score-matching network. Differently, our model is a conditional denoising diffusion probabilistic model which relies on a transformer-based architecture. Other diffusion-based approaches for time series (Li et al., 2022; Rasul et al., 2021; Tashiro et al., 2021) address slightly different problems, such as forecasting and imputation.
Market simulation with deep learning. Generating realistic market simulations using deep learning is a new paradigm. Traditional computational statistical approaches (Cont, 2001; Huang et al., 2015) and IABS (Byrd et al., 2020; Paddrik et al., 2012) rely on strong assumptions, such as constant order size, failing to produce realistic simulations. These methods are mainly used to study how the interactions of autonomous agents give rise to aggregate statistics and emergent phenomena in a system (Korczak and Hemes, 2017). Limit order book simulations are increasingly relying on deep learning. Li et al. (Li et al., 2020) were the first to leverage a Wasserstein GAN (WGAN) (Arjovsky et al., 2017) for generating order flows based on historical market data. Similarly, Coletta et al. (Coletta et al., 2021, 2022) employ WGANs in their stock market simulations, addressing the issue of responsiveness to experimental agents for the first time. Differently from (Coletta et al., 2021, 2022; Li et al., 2020), we condition with both the last orders and LOB snapshots, pushing the generation process toward a more realistic market simulation. Hultin et al. (Hultin et al., 2023) extend (Cont et al., 2010) and model individual features with separate conditional probabilities using RNNs. Instead of relying on GANs, which are prone to model collapse (Xiao et al., 2021) and instability (Chu et al., 2020), and RNNs, often hampered by the vanishing gradient phenomenon, we exploit diffusion-based models with an underlying transformer architecture. Nagy et al. (Nagy et al., 2023) rely on simplified state-space models (Smith et al., 2023) to learn long-range dependencies, tackling the problem via a masked language modeling approach (Devlin et al., 2018). Shi and Cartlidge (Shi and Cartlidge, 2023) introduce NS-LOB, a novel hybrid approach that combines a pre-trained neural Hawkes (Shi and Cartlidge, 2022) process with a multi-agent trader simulation. We refer the reader to (Jain et al., 2024) for a comprehensive review of limit order book simulations.
@@ 3. Related Works
Figure 1. The training procedure and architecture of TRADES. We apply noise to the time series’s last element until obtaining $x_{T}^{g}$. We condition $x_{T}^{g}$ on the previous $N-1$ orders by concatenating it to them on the sequence axis and the last $N$ LOB snapshots. We feed these new tensors to two separate MLPs, each composed of two fully connected layers, to augment them into a higher dimensional space. We concatenate the augmented output vectors on the features axis and sum the diffusion embedding step $t$ and the positional embedding; we detail how they are represented in the Appendix 4. We feed the result to the TRADES modules, composed of multi-head self-attention and feedforward layers, producing the noise $\varepsilon_{\theta}$ and the standard deviation $\Sigma_{\theta}$. $\varepsilon_{\theta}$ and $\Sigma_{\theta}$ go through a de-augmentation phase via MLPs to map them back to the input space and are used to reconstruct $\mathbf{x}_{T-1}^{g}$. We repeat this procedure $T$ times until we recover the original $\mathbf{x}_{0}^{g}$.
4. Transformer-based Denoising Diffusion Probabilistic Engine for LOB Simulations
@@
4.1. Generation with TRADES
Here, we focus on an abstract time series generation task with TRADES. The goal of probabilistic generation is to approximate the true conditional data distribution $q(\mathbf{x}_{0}^{g}\,|\,\mathbf{x}_{0}^{c})$ with a model distribution $p_{\theta}(\mathbf{x}_{0}^{g}\,|\,\mathbf{x}_{0}^{c})$. During the forward process, we apply noise only to the “future” – i.e., the part of the input we want to generate – while keeping the observed values unchanged. Therefore, the forward process is defined as in the unconditional case in Eq. (11) (Appendix). For the reverse process, we extend the unconditional one $p_{\theta}(\mathbf{x}_{0:T})$, defined in Eq. (12) (Appendix), to the conditional case:
(2) $p_{\theta}(\mathbf{x}_{0:T}^{g}):=p(\mathbf{x}_{T}^{g})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}^{g}\,|\,\mathbf{x}_{t}^{g},\mathbf{x}^{c}_{0})$
(3) $p_{\theta}(\mathbf{x}_{t-1}^{g}\,|\,\mathbf{x}_{t}^{g},\mathbf{x}^{c}_{0}):=\mathbf{\mathcal{N}}(\mathbf{x}_{t-1}^{g};\>\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t}^{g},\mathbf{x}^{c}_{0},t),\mathbf{\Sigma}_{\theta}(\mathbf{x}_{t}^{g},\mathbf{x}^{c}_{0},t))$

### v3

@@ Abstract
Financial markets are complex systems characterized by high statistical noise, nonlinearity, volatility, and constant evolution. Thus, modeling them is extremely hard. Here, we address the task of generating realistic and responsive Limit Order Book (LOB) market simulations, which are fundamental for calibrating and testing trading strategies, performing market impact experiments, and generating synthetic market data. We propose a novel TRAnsformer-based Denoising Diffusion Probabilistic Engine for LOB Simulations (TRADES). TRADES generates realistic order flows as time series conditioned on the state of the market, leveraging a transformer-based architecture that captures the temporal and spatial characteristics of high-frequency market data. There is a notable absence of quantitative metrics for evaluating generative market simulation models in the literature. To tackle this problem, we adapt the predictive score, a metric measured as an MAE, to market data by training a stock price predictive model on synthetic data and testing it on real data. We compare TRADES with previous works on two stocks, reporting a $\times 3.27$ and $\times 3.48$ improvement over SoTA according to the predictive score, demonstrating that we generate useful synthetic market data for financial downstream tasks. Furthermore, we assess TRADES’s market simulation realism and responsiveness, showing that it effectively learns the conditional data distribution and successfully reacts to an experimental agent, giving sprout to possible calibrations and evaluations of trading strategies and market impact experiments. To perform the experiments, we developed DeepMarket, the first open-source Python framework for LOB market simulation with deep learning. In our repository, we include a synthetic LOB dataset composed of TRADES’s generated simulations.
[…]
5552
1 Introduction
@@
A realistic and responsive market simulation [^1] To abbreviate, throughout this paper, by “market simulation”, we refer to a Limit Order Book (LOB) market simulation for a single stock. has always been a dream in the finance world [raberto2005modeling, levy2000microscopic, mizuta2016brief, jacobs2004financial]. Recent years have witnessed a surge in interest towards deep learning-based market simulations [li2020generating, nagy2023generative, cont2023limit, coletta2021towards]. An ideal market simulation should fulfill four key objectives: (1) enable the calibration and evaluation of algorithmic trading strategies including reinforcement learning models; (2) facilitate counterfactual experiments to analyze (2.1) the impact of orders [webster2023handbook], and (2.2) the consequences of changing financial regulations and trading rules, such as price variation limits; (3) analyze market statistical behavior and stylized facts in a controlled environment; (4) generate useful granular market data to facilitate research on finance and foster collaboration.
@@
For these objectives, two key elements are paramount: the realism and the responsiveness of the simulation. Realism refers to the similarity between the generated probability distribution and the actual data distribution. Responsiveness captures how the simulated market reacts to an experimental agent’s actions. Furthermore, the generated data usefulness is crucial for achieving the last objective (4). Usefulness refers to the degree to which the generated market data can contribute to other related financial tasks, such as predicting the trends of stock prices [prata2023lobbased, sirignano2019deep].
@@
Backtesting and Interactive Agent-Based Simulations (IABS) [byrd2020abides] are two of the most used traditional market simulation methods. Backtesting assesses the effectiveness of trading strategies on historical market data. It is inherently non-responsive since there is no way to measure the market impact of the considered trading strategies, making the analysis partial. IABS, on the other hand, enables the creation of heterogeneous pools of simplified traders with different strategies, aiming to approximate the diversity of the real market. However, obtaining realistic multi-agent simulations is challenging, as individual-level historical data of market agents is private, which impedes the calibration of the trader agents, resulting in an oversimplification of real market dynamics.
@@
Recent advances in generative models, particularly Wasserstein GANs [coletta2022learning, coletta2021towards, li2020generating], have shown promise in generating plausible market orders for simulations. However, GANs are susceptible to mode collapse [thanh2020catastrophic] and training instability [chu2020smoothness], leading to a lack of realism and usefulness in the generated data.
@@
To address these shortcomings, we present our novel TRAnsformer-based Denoising Diffusion Probabilistic Engine for LOB market Simulations (TRADES). TRADES generates realistic order flows that are time series, conditioned on past observations. We demonstrate that TRADES surpasses state-of-the-art (SoTA) methods in generating realistic and responsive market simulations. Importantly, due to its ability to handle multivariate time series generation, TRADES is adaptable to other domains requiring conditioned sequence data generation. Furthermore, TRADES readily adapts to an experimental agent introduced into the simulation, facilitating counterfactual market impact experiments. In summary, our contributions are:
  1. 1.
@@
    Realistic and responsive market simulation method: We develop a Denoising Diffusion Probabilistic Engine for LOB Simulations (TRADES), exploiting a transformer-based neural network architecture.
  2. 2.
@@
    DeepMarket: We release DeepMarket, the first open-source Python framework for market simulation with deep learning. We also publish TRADES’s implementation and checkpoints to promote reproducibility and facilitate comparisons and further research.
  3. 3.
@@
  4. 4.
    New “world” agent for market simulations: We extend ABIDES [byrd2020abides], an agent-based simulation environment, in our framework by introducing a new world agent class accompanied by a simulation configuration, which, given in input a trained generative model, creates limit order book market simulations. Our experimental framework does not limit the simulation to a single-world agent but enables the introduction of other trading agents, which interact among themselves and with the world agent. This defines a hybrid approach between deep learning-based and agent-based simulations.
  5. 5.
    First quantitative metric for market simulations: The literature shows a notable absence of quantitative metrics to evaluate the generated market simulations. Typically, the evaluation relies on plots and figures. We posit that a robust and quantitative assessment is essential for the comparative analysis of various methodologies. To this end, we adapt the predictive score introduced in [yoon2019time] to objectively and quantitatively evaluate the usefulness of the generated market data.
  6. 6.
    Extensive experiments assessing usefulness, realism, and responsiveness: We perform a suite of experiments to demonstrate that TRADES-generated market simulations abide by these three principles. We show how TRADES outperforms SoTA methods [byrd2020abides, coletta2022learning, coletta2021towards, vyetrenko2020get] according to the adopted predictive score and illustrate how TRADES follows the established stylized facts in financial markets [vyetrenko2020get].
2 Background
Here, we provide background information on multivariate time series generation and limit order book markets. Furthermore, since TRADES is an extension of the Denoising Diffusion Probabilistic Model (DDPM), we summarize it in the Appendix [bertiecaiappendix].
2.1 Multivariate time series generation
@@
$q(\mathbf{x})=\prod_{n=1}^{N}q(\mathbf{x}_{n}\>|\>\mathbf{x}_{1},...,\,\mathbf{x}_{n-1})$ (1)
@@
2.2 Limit Order Book
@@
In a Limit Order Book (LOB) market, traders can submit orders to buy or sell a certain quantity of an asset at a particular price. There are three main types of orders in a limit order market. (1) A market order is filled instantly at the best available price. (2) A limit order allows traders to decide the maximum (in the case of a buy) or the minimum (in the case of a sell) price at which they want to complete the transaction. A quantity is always associated with the price for both types of orders. (3) A cancel order [^4] Sometimes, events referring to these orders are defined as deletion. removes an active limit order. The Limit Order Book (LOB) is a data structure that stores and matches the active limit orders according to a set of rules. The LOB is accessible to all the market agents and is updated with each event, such as order insertion, modification, cancellation, and execution. The most used mechanism for matching orders is the Continuous Double Auction (CDA). In a CDA, orders are executed whenever a price overlaps between the best bid (the highest price a buyer is willing to pay) and the best ask (the lowest price a seller is willing to accept). This mechanism allows traders to trade continuously and competitively. The evolution over time of a LOB represents a multivariate temporal problem. We can classify the research on LOB data into four main types of studies, namely empirical studies analyzing the LOB dynamics [bouchaud2002statistical, cont2001empirical], price and volatility forecasting [zhang2019deeplob, sirignano2019deep], modeling the LOB dynamics [cont2011statistical, gould2013limit] and LOB market simulation [byrd2020abides, coletta2021towards, li2020generating].
@@
Figure 1: The training procedure and architecture of TRADES. We apply noise to the time series’s last element until obtaining $x_{T}^{g}$. We condition $x_{T}^{g}$ on the previous N1N-1 orders and the last $N$ LOB snapshots. We feed these new tensors to two separate MLPs, each composed of two fully connected layers, to augment them into a higher-dimensional space. We concatenate the augmented output vectors on the features axis and sum the diffusion embedding step $t$ and the positional embedding; We feed the result to the TRADES modules, composed of multi-head self-attention and feedforward layers, producing the noise $\varepsilon_{\theta}$ and the standard deviation $\Sigma_{\theta}$. $\varepsilon_{\theta}$ and $\Sigma_{\theta}$ go through a de-augmentation phase via MLPs to map them back to the input space and are used to reconstruct $\mathbf{x}_{T-1}^{g}$. We repeat this procedure $T$ times until we recover the original $\mathbf{x}_{0}^{g}$.
3 Related Works
Diffusion models for time series generation. Diffusion models have been successfully applied to generate images [austin2021structured], video [mei2023vidm], and text [zhu2023conditional]. Recently, they have also been exploited for time series forecasting [li2023generative], imputation [tashiro2021csdi], and generation [lim2023regular]. Lim et al. [lim2023regular] tackle time series generation using diffusion models. They present TSGM, which relies on an RNN-based encoder-decoder architecture with a conditional score-matching network. Yangming Li’s TS-Diffusion [yuan2024diffusionts] introduces a neural-ODE encoder with self-attention and a diffusion core to manage irregularities and missingness in complex series
Market simulation with deep learning. Generating realistic market simulations using deep learning is a new paradigm. Traditional computational statistical approaches [cont2001empirical] and IABS [byrd2020abides] rely on strong assumptions, such as constant order size, failing to produce realistic simulations. These methods are mainly used to study how the interactions of autonomous agents give rise to aggregate statistics and emergent phenomena in a system. Limit order book simulations are increasingly relying on deep learning. Li et al. [li2020generating] were the first to leverage a Wasserstein GAN (WGAN) [arjovsky2017wasserstein] for generating order flows based on historical market data. Similarly, Coletta et al. [coletta2021towards, coletta2022learning] employ WGANs in their stock market simulations, addressing the issue of responsiveness to experimental agents for the first time. Different
from [coletta2021towards, coletta2022learning, li2020generating], we condition with both the last orders and LOB snapshots, pushing the generation process toward a more realistic market simulation. Hultin et al. [hultin2023generative] extend [cont2010stochastic] and model individual features with separate conditional probabilities using RNNs. Instead of relying on GANs, which are prone to model collapse and instability [chu2020smoothness], and RNNs, often hampered by the vanishing gradient phenomenon, we exploit diffusion-based models with an underlying transformer architecture. Nagy et al. [nagy2023generative] rely on simplified state-space models to learn long-range dependencies, tackling the problem via a masked language modeling approach. Shi and Cartlidge [shi2023neural] introduce NS-LOB, a novel hybrid approach that combines a pre-trained neural Hawkes [shi2022state] process with a multi-agent trader simulation. We refer the reader to [jain2024limit] for a comprehensive review of limit order book simulations.
4 Transformer-based Denoising Diffusion Probabilistic Engine for LOB Simulations
@@
4.1 Generation with TRADES
Here, we focus on an abstract time series generation task with TRADES. The goal of probabilistic generation is to approximate the true conditional data distribution $q(\mathbf{x}_{0}^{g}\,|\,\mathbf{x}_{0}^{c})$ with a model distribution $p_{\theta}(\mathbf{x}_{0}^{g}\,|\,\mathbf{x}_{0}^{c})$. During the forward process, we apply noise only to the “future” – i.e., the part of the input we want to generate – while keeping the observed values unchanged. Therefore, the forward process is defined as in the unconditional case (see Appendix [bertiecaiappendix] or [ho2020denoising]). For the reverse process, we extend the unconditional
one $p_{\theta}(\mathbf{x}_{0:T})$, to the conditional case:
$p_{\theta}(\mathbf{x}_{0:T}^{g}):=p(\mathbf{x}_{T}^{g})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}^{g}\,|\,\mathbf{x}_{t}^{g},\mathbf{x}^{c}_{0})$ (2)
$p_{\theta}(\mathbf{x}_{t-1}^{g}\,|\,\mathbf{x}_{t}^{g},\mathbf{x}^{c}_{0}):=\mathbf{\mathcal{N}}(\mathbf{x}_{t-1}^{g};\>\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t}^{g},\mathbf{x}^{c}_{0},t),\mathbf{\Sigma}_{\theta}(\mathbf{x}_{t}^{g},\mathbf{x}^{c}_{0},t))$ (3)